OpenStudy (anonymous):

A proportional relationship between the number of pounds of potatoes (x) and the price in dollars (y) is graphed, and the ordered pair (4, 3) is on the graphed line. Part A: What is the price of 1 pound of potatoes? Show your work, including the proportion you used to determine the price. (8 points) Part B: What does the ordered pair (8, 6) on the graph represent? Explain in words.

OpenStudy (anonymous):

Morning becky

OpenStudy (anonymous):

Part A: I know I have 3 = 4m + b, but I cant solve yet since I dont have two variables. So we have 3 = 3/4 x 4 + b. The answer would be 0. So y is 0. Now we have the equation y = 3/4x +0. Y is = to 3/4. i wrote this for part a.

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

umm. Well first morning @isaac4321 second wdym @adcooks

OpenStudy (anonymous):

umm. Well first morning @isaac4321 second wdym @adcooks

OpenStudy (anonymous):

I don't really know what to do I just wanted to say heyyyyy

OpenStudy (anonymous):

...Really lol